OpenStudy (swaugh):

FUN MATH GAME Match each equation with the ordered pair that is a solution. 1.y=2x-1 A.(1,1) 2.y=2x-6 B.(2, 4) 3.y=x+3 C.(3, 6) 4.y=3x-2 D.(5, 4)

ganeshie8 (ganeshie8):

A.(1,1) plug this in first equation and see if it satisfies

ganeshie8 (ganeshie8):

y = 2x - 1 plugin x = 1, y = 1 above. what do you get ?

OpenStudy (swaugh):

1

ganeshie8 (ganeshie8):

Yes : y = 2x - 1 1 = 2*1 - 1 1 = 2 - 1 1 = 1 this is true, therefore 1st equation matches with A

OpenStudy (swaugh):

thats how you do it

OpenStudy (swaugh):

you plug in the numbers

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

Yes, plugin B(2, 4) in the remaining equations and see which one satisfies

OpenStudy (swaugh):

okay

OpenStudy (swaugh):

2 is B

ganeshie8 (ganeshie8):

Nope, try again

OpenStudy (mathstudent55):

Is this a novel way of getting someone to do your homework/test/quiz, to call is a "fun math game"? That's brilliant.

OpenStudy (swaugh):

it has to be B

OpenStudy (swaugh):

2 = B